OpenStudy (anonymous):

How to tabulate based on the graph that being produced by mathlab.? This is my program

OpenStudy (anonymous):

L = 1.; T = 1.; maxk = 2500; dt = T/maxk; n=50; dx = L/n; cond = 1/4; b = 2.*cond*dt/(dx*dx); for i= 1:n+1 x(i) = (i-1)*dx; u(i,1) = sin(pi*x(i)); end for k=1:maxk+1 u(1,k) = 0.; u(n+1,k)=0.; time(k) = (k-1)*dt; end for k=1:maxk for i=2:n; u(i,k+1)=u(i,k)+0.5*b*(u(i-1,k)+u(i+1,k)-2.*u(i,k)); end end figure (1) plot (x,u(:,1),'-',x,u(:,100),'-',x,u(:,300),'-',x,u(:,600),'-') title ('Temperature within the explicit method') xlabel ('X') ylabel ('Y') figure (2) mesh (x,time,u') title ('Temperature within the explicit method') xlabel ('X') ylabel ('Temperature')

OpenStudy (anonymous):

I need to tabulate the data for u(x,t) vs x. Someone help me please :( this is the program.